# Evaluating Perspectival Biases in Cross-Modal Retrieval

### v1

@@ Abstract
Multimodal retrieval systems are expected to operate in a semantic space, agnostic to the language or cultural origin of the query. In practice, however, retrieval outcomes systematically reflect perspectival biases: deviations shaped by linguistic prevalence and cultural associations. We study two such biases. First, prevalence bias refers to the tendency to favor entries from prevalent languages over semantically faithful entries in image-to-text retrieval. Second, association bias refers to the tendency to favor images culturally associated with the query over semantically correct ones in text-to-image retrieval. Results show that explicit alignment is a more effective strategy for mitigating prevalence bias. However, association bias remains a distinct and more challenging problem. These findings suggest that achieving truly equitable multimodal systems requires targeted strategies beyond simple data scaling and that bias arising from cultural association may be treated as a more challenging problem than one arising from linguistic prevalence.
Keywords model bias/fairness evaluation $\cdot$ multimodality $\cdot$ multilingual evaluation $\cdot$ language/cultural bias analysis
@@ 1 Introduction
As Nietzsche [1] observed, “there is only a perspective seeing, only a perspective knowing”; put differently, there is no view from nowhere. Large models inherit this perspectival character through their training data; what they learn to represent depends on the frequency of appearance and co-occurrence structure. As a result, the latent space of such models does not always function as the robust, language-agnostic semantic space we expect. Instead, retrieval outcomes can be skewed, favoring linguistic prevalence or cultural association over true semantic relevance. The effect of such a perspectival character on both image-to-text and text-to-image retrievals is illustrated in Figure 1. Understanding and quantifying these effects is crucial for ensuring consistent retrieval performance across languages and cultures.
@@
Figure 1: Two Forms of Perspectival Biases. (a) Prevalence bias: an image query favors high-resource languages. A retrieval model places English results above semantically equivalent Japanese and Thai captions. (b) Association Bias: A visualized model’s embedding space, demonstrating how a Japanese text query for “necklace” retrieves culturally proximate images (Japanese masks) instead of the semantically correct one (Kenyan necklace).
Multimodal retrieval enables cross-modality search, primarily between text and images. Early models, such as CLIP [2], align vision and language representations through paired supervision. Recent Multimodal Large Language Models (MLLMs) [3, 4] achieve alignment implicitly through large-scale pretraining. Despite these advancements, the critical issue of language and cultural bias in retrieval remains underexplored.
This lack of study is concerning given that state-of-the-art retrievers are trained on web-scale, text-image datasets like LAION [5] and WebLi [6], which are overwhelmingly English-centric. While these datasets are constructed using English alt-text, images with high cultural specificity often retain alt-text in their native languages. As observed in multilingual food datasets [7], items like the Catalan pastry “coca de recapte” are exclusively described in Catalan or Spanish. This might allow models to develop emergent multilingual capabilities, but it also risks introducing a systemic bias where the model learns spurious correlations, preferentially matching images with text from a specific majority or “expected” language.
A key barrier to investigating these biases is the absence of targeted metrics and benchmarks designed to quantify them. To address this gap, we introduce an evaluation framework for both retrieval directions, each capturing a distinct form of perspectival bias. Image-to-text retrieval. In the absence of linguistic cues, retrievals reveal how the prevalence of certain languages in the training data shapes the results. To assess this prevalence bias, we propose the Discounted Language Bias Kullback–Leibler Divergence (DLBKL), inspired by Language Bias Kullback–Leibler Divergence (LBKL) [8], which measures how strongly retrieval relevance depends on language rather than semantics, as shown in Figure 1a. Text-to-image retrieval. When linguistic and cultural cues are present in the query, retrievals reveal the model’s tendency to favor culturally associated visual patterns over semantically aligned ones. We term this association bias and construct a balanced, cross-cultural, and cross-lingual dataset to disentangle semantic relevance from cultural proximity, as shown in Figure 1b.
Using these tools, we conduct an empirical analysis comparing the perspectival biases inherent in retrievers adapted from MLLMs with those trained using explicit cross-lingual alignment techniques [9, 10]. Our findings reveal that models with explicit alignment mechanisms exhibit lower biases, highlighting a critical trade-off between the scale of MLLMs and the fairness of more targeted alignment strategies.
We summarize our contributions as follows: (i) We propose DLBKL, a metric for quantifying prevalence bias in multimodal retrieval within a multilingual candidate pool to assess language fairness. (ii) We introduce a novel vision-language dataset, parallel across culture and language, designed to assess association bias.
2 Related work
2.1 Multimodal Retrievers
Retrieving image information using a text query can be accomplished by two methods: sparse and dense retrieval. Sparse retrieval methods utilize a high-dimensional representation extracted from words in a text query or an image caption [11, 12]. While these methods are fast, they do not understand the semantics of the image as they solely rely on the image caption to represent the content. To handle the challenge of understanding semantic meaning, deep learning-based dense retrieval methods have been developed. For example, CLIP [2] and ALIGN [13] used a dual-encoder architecture specifically trained to connect the semantic meanings of text-image pairs using contrastive loss objectives. Both models were built on similar principles but varied in text-image data size and utilized different text/image encoders. The recent ColQwen [14] and GME [15] model adapts a Large Language Model (LLM) to learn a more intensive semantic connection between text and images, converting them to a multimodal retrieval model.
2.2 Language Bias in Multi-model Retrievers
Language bias in multimodal retrieval refers to performance differences that arise when semantically equivalent queries in different languages yield divergent rankings, often favoring high-resource languages such as English. Prior work frames this along two fairness axes: (i) an individual-level notion, where multilingual variants of the same query should produce similar results, and (ii) a group-level notion, where aggregate retrieval performance should remain balanced across languages [16].
A study on multilingual retrieval benchmark [17] reports uneven performance across languages, with comparatively stronger results on English and other high-resource languages for various modern multimodal retrievers, despite their large scale in data and parameters. For instance, some models exhibit significant variation in NDCG [18] scores across different languages, indicating that retrieval effectiveness is not uniform.
To quantify such disparities, fairness-aware metrics from ranking literature, such as exposure parity [18], have been adapted to language as a protected attribute. More recently, Adewumi et al. [19] surveyed multimodal bias, emphasizing the lack of dedicated language-focused evaluation protocols. Addressing this, Laosaengpha et al. [8] proposed LBKL, a distributional measure of divergence between retrieval results across language variants. Although LBKL was designed for measuring text modality bias, it can technically be extended to multimodal retrieval, enabling a more fine-grained detection of scores across languages. These works highlight several metrics for measuring language bias in multimodal retrievers. However, despite their effectiveness in measuring language bias, none take retrieval rankings into account.
2.3 Multilingual and Cross-Lingual Retrieval Strategies
Two dominant paradigms exist for building multilingual multimodal retrieval systems: holistic end-to-end pre-training and explicit cross-lingual alignment.
Holistic End-to-End Pre-training on Large-Scale Multilingual Data
This approach, anchored by foundation MLLMs like Qwen2.5-VL [3], aims to learn an emergent universal representation from web-scale, mixed-language data. Within this paradigm, some models like GME [15] and the standard ColQwen series [14] are fine-tuned on predominantly English datasets. Others, such as the multilingual ColQwen series [14] and jina-embeddings-v4 [20], intentionally incorporate extensive multilingual data to improve fairness.
Explicit Cross-Lingual Alignment via Knowledge Distillation
This alternative strategy uses knowledge distillation to align text encoders for new languages to a strong, pre-existing English model’s embedding space, such as CLIP’s. This data-efficient method, exemplified by M-CLIP [10], typically requires only parallel text corpora to force non-English embeddings to mimic their English counterparts via a teacher-student setup.
Our work evaluates models representing both paradigms, providing a direct comparison of biases inherent to each approach.
3 Methodology for Evaluating Bias in Multimodal Retrieval
@@
This section outlines the framework developed to investigate perspectival bias in multilingual, multimodal retrieval systems. We first state our guiding research questions and then explain the studies that address these research questions in Sections 3.1 and 3.2. Our investigation centers on two complementary perspectives, corresponding to different retrieval directions, as shown in Figure 2:
@@ 2. RQ2
Figure 2: Overview of the study. First, in RQ1, we identify language prevalence in image-to-text retrieval by analyzing the language of the retrieved text and comparing it with high-resource languages, as well as medium- and low-resource languages. Second, in RQ2, we identify the association bias using self-preference behavior of the model by retrieving an image with three candidates: semantically relevant, culturally relevant, and a non-relevant candidate.
3.1 RQ1: Image-to-Text Retrieval Study
In this study, we assess the bias arising from linguistic prevalence by examining the discrepancy between an expected “fair” language distribution and the observed one. That is, the discrepancy should be zero if the linguistic prevalence has no effect on the retrieval results and increases as the results deviate from the ideal case. As discussed in Section 2.2, existing work lacks a dedicated metric to quantify such a discrepancy in multimodal retrieval.
As the first step to closing this gap, we apply the Language Bias Kullback–Leibler (LBKL) Divergence proposed by Laosaengpha et al. [8], which measures the divergence between an expected language distribution and the observed distribution in a retrieved list. Given proportions of language A and B in the ground truth ($P_{A}(x),P_{B}(x)$) and in the retrieved set ($Q_{A}(x),Q_{B}(x)$), LBKL is given as:
@@ 3.1 RQ1: Image-to-Text Retrieval Study
While LBKL can be applied to cross-modal retrieval, it is rank-agnostic: deviations at rank 1 are penalized equally to deviations at rank 100. This underestimates the harm in systems that concentrate resource-driven bias near the top ranks.
For the next step, we extend LBKL by introducing the Discounted Language Bias Kullback-Leibler (DLBKL) divergence, which incorporates a logarithmic rank discount inspired by NDCG [18]. We assign a weight $w(i)=1/\log_{2}(i+1)$ to each rank $i$. The rank-weighted proportion for a language $l$ is then:
@@
where $\mathbb{I}(\cdot)$ is the indicator function. DLBKL is calculated by substituting $Q^{\prime}_{l}(x)$ for $Q_{l}(x)$ in the LBKL formula. As illustrated in Figure 3, DLBKL penalizes top-ranked disparities more heavily, aligning the metric with user exposure and better capturing the discrepancy between the ideal case and observed one in multimodal retrieval.
Figure 3: Illustration of how DLBKL, unlike the rank-agnostic LBKL, assigns a higher bias score to lists where high-resource languages dominate the top ranks.
3.2 RQ2: Text-to-Image Retrieval Study
@@
To quantify the degree to which models prioritize cultural association over semantic fidelity, a phenomenon we illustrate in Figure 1(b), a benchmark with a parallel structure in its cultural dimension is necessary. To the best of our knowledge, no such benchmark exists, so we make two primary contributions. First, we construct and introduce the Cross-Cultural Multimodal (3XCM) benchmark, a novel dataset designed specifically for this purpose. Second, we propose the Self-Preference Cultural Bias Score (SP), a new metric for explicitly measuring this form of bias.
3.2.1 The 3XCM Dataset Benchmark
To evaluate association bias, we constructed the 3XCM benchmark [^1] Research release only (CC BY-NC-SA 4.0). Ethical review required for production use.
Available at: https://huggingface.co/datasets/Chula-AI/association_bias_benchmark.
. The process involved two primary stages: (i) gathering a corpus of culturally diverse images and (ii) structuring these images into a triplet-based evaluation set.
@@ 3.2.1 The 3XCM Dataset Benchmark
  • Concept Generation We used Gemini [^2] Version used: gemini-2.5-flash (Released June 17, 2025). to generate a large pool of concepts, which we manually curated to a final set of 138 coarse-grained, culturally-inclusive concepts (e.g., "train", "food"). Each concept is an abstract, semantic category that uses shared properties to group a broad, culturally-inclusive range of entities. The prompt for generating concepts can be found in Appendix A.
@@
  • Concept De-duplication: We use BGE-M3 [21] to de-duplicate concepts based on similarity with a threshold of 0.92.
@@
  • Image Collection: For each concept and a set of 16 diverse countries, we used the DuckDuckGo image search API [22] to retrieve the top 10 images using queries in both English (e.g., "train Japan") and the local native language.
@@
  • Image De-duplication: To ensure visual diversity, we performed two-stage de-duplication within each concept. First, near-exact duplicates were removed automatically using an embedding model. Subsequently, three human annotators, following the guidelines in Appendix F, used a custom tool to manually filter out remaining images that depicted the same scene or object without meaningful variation in viewpoint or time of day.
Leveraging the collected cultural images, we introduce a novel evaluation paradigm that employs a forced-choice task. This setup is designed to disambiguate between the model’s reliance on semantic understanding (the concept) and its preference for cultural association. As illustrated in Figure 5, for a given query (e.g., "food" in Thai), the model is presented with a triplet of image candidates: (i) Semantically Relevant: same concept, different culture (e.g., Nigerian food); (ii) Culturally Relevant: different concept, same culture (e.g., Thai traditional dance); and (iii) Non-Relevant: different concept and culture (e.g., Japanese gas station).
Figure 5: Illustration of association bias evaluation. A Thai text query for “food” is evaluated against three candidates designed to isolate semantic faithfulness vs. cultural relevance.
@@
The final dataset contains 11,724 entries distributed across 138 concepts. Further statistics and samples are provided in Appendix J and N respectively.
3.2.2 Self-Preference Cultural Bias Score (SP)
@@
With the constructed dataset, we can now measure the discrepancy between the ideal case and the observed one, where bias arising from cultural association may intervene. Ideally, the discrepancy should be zero when image retrieval depends solely on semantic relevance, and it should increase as the model’s preference tends towards images associated with the culture of the query, rather than semantic accuracy. To quantify the discrepancy, we propose a metric called the self-preference cultural bias score (SP), which can be computed as follows:
$M_{k}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(S_{k,i}=\max(S_{\text{sem},i},S_{\text{cul},i},S_{\text{non},i})\right)$ (3)
$\text{SP}=\frac{M_{\text{cul}}}{M_{\text{sem}}}$ (4)
@@
where $M_{k}$ is the proportion of times a candidate of type $k$ receives the highest similarity score across $N$ total trials. The candidate type $k$ can be semantically relevant (sem), culturally relevant (cul), or non-relevant (non). The similarity score for candidate type $k$ in trial $i$ is denoted by $S_{k,i}$. The indicator function $\mathbb{I}(\cdot)$ is 1 if the condition is true and 0 otherwise. The SP score (Eq. 4) is then the ratio of cultural wins ($M_{\text{cul}}$) to semantic wins ($M_{\text{sem}}$). In this way, a higher SP score indicates stronger cultural self-preference over semantic faithfulness and thus a greater extent of association bias.

### v3

@@ Abstract
Multimodal retrieval systems are expected to operate in a semantic space, agnostic to the language or cultural origin of the query. In practice, however, retrieval outcomes systematically reflect perspectival biases: deviations shaped by linguistic prevalence and cultural associations. We introduce the Cross-Cultural, Cross-Modal, Cross-lingual Multimodal (3XCM) benchmark to isolate these effects. Results from our studies indicate that, for image-to-text retrieval, models tend to favor entries from prevalent languages over those that are semantically faithful. For text-to-image retrieval, we observe a consistent “tugging effect” in the joint embedding space between semantic alignment and language-conditioned cultural association. When semantic representations are insufficiently resolved, particularly in low-resource languages, similarity is increasingly governed by culturally familiar visual patterns, leading to systematic association bias in retrieval. Our findings suggest that achieving equitable multimodal retrieval necessitates targeted strategies that explicitly decouple language from culture, rather than relying solely on broader data exposure. This work highlights the need to treat linguistic and cultural biases as distinct, measurable challenges in multimodal representation learning.
@@ 1 Introduction
As nietzsche1887genealogy observed, there is only a perspective seeing, only a perspective knowing”; put differently, there is no view from nowhere. Large models inherit this perspectival character from training data, where representations depend on frequency and co-occurrence. Consequently, their latent spaces deviate from the expected robust, language-agnostic semantics, skewing retrieval toward linguistic prevalence or cultural associations over true relevance. Figure 1 illustrates this in image-to-text and text-to-image retrieval, highlighting the need to quantify such effects for consistent cross-lingual, cross-cultural performance.
@@
Figure 1: Two Forms of Perspectival Biases. (a) Prevalence bias: an image query favors high-resource languages. A model places English results above semantically equivalent Japanese and Thai captions. (b) Association Bias: Given a Japanese text query for “necklace”, a model places a culturally proximate image (Japanese masks) closer to the query than a semantically correct one (Kenyan necklace).
Multimodal retrieval aligns text and images via paired supervision in early models like CLIP [DBLP:journals/corr/abs-2103-00020] or implicit pretraining in recent Multimodal Large Language Models (MLLMs) [bai2025qwen25vltechnicalreport, comanici2025gemini]. Yet, language and cultural biases remain underexplored, especially given English-centric datasets like LAION [schuhmann2022laion5bopenlargescaledataset] and WebLi [chen2023pali]. While culturally specific images may carry native-language alt-text, e.g., Catalan descriptions for “coca de recapte” [olondriz2021foodi]. This fosters emergent multilingualism but risks spurious correlations, favoring “expected” languages.
A major barrier is the lack of targeted benchmarks and metrics. To address this, we develop an evaluation framework isolating two perspectival biases across retrieval directions. For image-to-text (lacking linguistic cues), we assess prevalence bias via the proposed Discounted Language Bias Kullback–Leibler Divergence (DLBKL), extending LBKL [laosaengpha2025mitigatinglanguagebiascrosslingual] to penalize high-resource languages dominating top ranks (Figure 1a). For text-to-image (with linguistic/cultural cues), we measure association bias using a parallel, cross-cultural dataset to disentangle semantic fidelity from cultural proximity (Figure 1b). Extending this, we test the ability of explicit cultural descriptors (CDs) to override association bias, examining the “tugging effect” where models struggle to prioritize CDs (e.g., “Japanese train” in Thai) over implicit query-language associations.
Using these tools, we compare biases in MLLM-adapted retrievers against those with explicit cross-lingual alignment [chen-etal-2023-mclip, carlsson-etal-2022-cross]. Findings expose biases in both directions, including CDs’ limited efficacy: models follow instructions but fall back to language-linked visuals when cultural semantics are unresolved.
Our contributions: (i) DLBKL as a supplementary, rank-sensitive extension to LBKL for measuring prevalence bias in multilingual pools; (ii) the 3XCM benchmark, parallel across cultures/languages for association bias; (iii) analysis of CDs’ impact on retrieval, revealing persistent biases despite explicit guidance.
2 Related Works
2.1 Language Bias in Multimodal Retrievers
Language bias in multimodal retrieval refers to performance disparities where high-resource languages dominate rankings for semantically equivalent queries. osmulski2025miracl report that modern retrievers exhibit significant variation in NDCG scores [ndcg] across languages, indicating that their effectiveness depends on resources and script types.
To quantify these disparities, fairness-aware metrics like exposure parity [ndcg] have been adapted, yet adewumi2024fairnessbiasmultimodalai emphasizes the lack of dedicated language-focused protocols. laosaengpha2025mitigatinglanguagebiascrosslingual introduced LBKL, a distributional measure of divergence for text modality bias. While technically extensible to multimodal retrieval, such metrics ignore retrieval rankings in the bias measurement.
2.2 Cultural Benchmarks
While recent benchmarks [liu-etal-2021-visually, romero2024cvqa, nayak-etal-2024-benchmarking, li2025ravenea] have made progress in evaluating cultural reasoning and knowledge, they primarily focus on culture-specific concepts or broad facets. In contrast, the assessment of bias discussed in Section 2.1 requires a basic-level, parallel corpus of universal concepts to strictly disentangle semantic relevance from cultural association in retrieval rankings. We provide a detailed comparison to other datasets in Appendix A.
3 Methodology
@@
This section outlines the framework developed to investigate perspectival bias in multilingual, multimodal retrieval systems. We first state our guiding research questions and subsequently detail the experiments addressing them in Sections 3.1, 3.2.2, and 3.2.3. Our investigation is organized into three specific research questions as follows:
@@ 2. RQ2
  3. RQ3
[…]
3.1 Image-to-Text Retrieval (RQ1)
To assess prevalence bias, we measure the divergence between expected and observed language distributions. We adopt the Language Bias Kullback–Leibler (LBKL) divergence [laosaengpha2025mitigatinglanguagebiascrosslingual]:
@@ 3.1 Image-to-Text Retrieval (RQ1)
Since LBKL is rank-agnostic, we propose Discounted LBKL (DLBKL) to prioritize penalizing bias in higher ranks more than lower ranks. We apply a logarithmic discount $w(i)=1/\log_{2}(i+1)$ to the observed language proportion $Q^{\prime}_{l}(x)$:
@@
DLBKL is computed by substituting $Q^{\prime}_{l}(x)$ for $Q_{l}(x)$ in equation 1. This formulation can be generalized to support cases where the bias assessment involves more than two languages by measuring divergence over the full distribution. It penalizes models where high-resource languages disproportionately occupy top ranks relative to the prior $P$. Crucially, in parallel datasets where semantic relevance is uniform across languages, this formulation isolates prevalence bias without compromising retrieval accuracy (see Figure 2 and Appendix D).
Figure 2: Illustration of how DLBKL, unlike the rank-agnostic LBKL, assigns a higher bias score to lists in which high-resource languages dominate the top ranks, under the setup that the prior P is a uniform distribution
3.2 Text-to-Image Retrieval (RQ2 & RQ3)
To quantify the degree to which models prioritize cultural association over semantic fidelity, a phenomenon we illustrate in Figure 1(b), a benchmark with a parallel structure in its cultural dimension is necessary. To the best of our knowledge, no such benchmark exists, so we make two primary contributions. First, we construct and introduce the Cross-Cultural, Cross-Modal, Cross-lingual Multimodal (3XCM) benchmark, a novel dataset designed specifically for this purpose. Second, we propose the Self-Preference Cultural Bias Score (SP), a new metric for explicitly measuring this form of bias.
3.2.1 The 3XCM Dataset Benchmark
@@
To evaluate association bias, we constructed the 3XCM benchmark, a challenge set designed to elicit culture-language association bias. This evaluation requires strict cultural parallelism at the concept level (e.g., "school" or "wedding" across all cultures), a structural feature that existing datasets [liu-etal-2021-visually, romero2024cvqa, nayak-etal-2024-benchmarking] do not provide. [^1] Research release only (CC BY-NC-SA 4.0). Ethical review required for production use. The process involved two primary stages: (i) gathering a corpus of culturally diverse images and (ii) structuring these images into a triplet-based evaluation set.
@@
  • Concept Generation We used Gemini [^2] Version used: gemini-2.5-flash (Released June 17, 2025). to generate a large pool of concepts, which we manually curated to a final set of 138 coarse-grained, culturally-inclusive concepts (e.g., "train", "food"). Each concept is an abstract, semantic category that uses shared properties to group a broad, culturally-inclusive range of entities. The prompt for generating concepts can be found in Appendix H.
@@
  • Concept De-duplication: We use BGE-M3 [chen2024bgem3embeddingmultilingualmultifunctionality] to de-duplicate concepts based on similarity with a threshold of 0.92.
@@
  • Image Collection: For each concept and a set of 16 diverse countries, we used the DuckDuckGo image search API [deepanprabhu:duckduckgo] to retrieve the top 10 images using queries in both English (e.g., "train Japan") and the local native language.
@@
  • Image De-duplication: To ensure visual diversity, we performed two-stage de-duplication within each concept. First, near-exact duplicates were removed automatically using an embedding model. Subsequently, three human annotators, following the guidelines in Appendix J, used a custom tool to manually filter out remaining images that depicted the same scene or object without meaningful variation in viewpoint or time of day.
@@
The final dataset contains 11,724 entries distributed across 138 concepts. Further statistics and samples are provided in Appendix K and L respectively.
3.2.2 RQ2: Effect of Implicit Association Bias
To evaluate implicit association bias, we set up a forced-choice task designed to disambiguate between the model’s reliance on semantic understanding (the concept) and its preference for cultural association. As illustrated in Figure 4(a), for a given query (e.g., "train" in Thai), the model is presented with a triplet of image candidates representing three answer categories: (i) Correct: The model successfully identifies an image with a semantically relevant object (e.g., a train of any culture). (ii) Language-biased: The model prioritizes culture-language association over semantic relevance (e.g., a Thai dance). (iii) Totally irrelevant: The model retrieves an image without any semantic or cultural relevance (e.g., the Eiffel Tower).
Figure 4: Illustration of association bias evaluation. (a) RQ2 evaluates a Thai query for “train” against three candidates to distinguish semantic faithfulness from cultural relevance. (b) RQ3 uses the query “train of Japan” to measure the tension between explicit cultural descriptors and implicit language-cultural associations.
Note that we deliberately exclude images with relevance in terms of object semantics and query language to expose the prioritization of relevance.
@@ 3.2.2 RQ2: Effect of Implicit Association Bias
With the constructed dataset, we can now measure the discrepancy between ideal retrieval outcomes and observed results, where bias arising from cultural association may intervene. Ideally, the discrepancy should be zero when image retrieval depends solely on semantic relevance, and it should increase as the model’s preference tends towards images associated with the culture of the query, rather than semantic accuracy. To quantify the discrepancy, we propose a metric called the self-preference cultural bias score (SP), which can be computed as follows:
$M_{k}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(S_{k,i}=\max(S_{\text{cr},i},S_{\text{lb},i},S_{\text{ti},i})\right)$ (3)
$\text{SP}=\frac{M_{\text{lb}}}{M_{\text{cr}}}$ (4)
@@
where $M_{k}$ is the proportion of times a candidate of type $k$ receives the highest similarity score across $N$ total trials. The candidate type $k$ can be Correct (cr), Language-biased (lb), or Totally Irrelevant (ti). The similarity score for candidate type $k$ in trial $i$ is denoted by $S_{k,i}$. The indicator function $\mathbb{I}(\cdot)$ is 1 if the condition is true and 0 otherwise. The SP score (equation 4) is then the ratio of language-biased wins ($M_{\text{lb}}$) to correct answer wins ($M_{\text{cr}}$). In this way, a higher SP score indicates stronger cultural self-preference over semantic faithfulness and thus a greater extent of association bias.
3.2.3 RQ3: Effect of Cultural Descriptors
To investigate the model’s adherence to explicit instructions against language-driven bias, we extend the forced-choice framework to include specific cultural cues. We construct queries using an “Object & Cultural Descriptor” format (e.g., “Train of Japan” formulated in Thai) to gauge the “tugging effect” between the explicit country mention and the implicit language-cultural association. As illustrated in Figure 4(b), we expand the answer category to six to isolate specific retrieval failures: (i) Correct: The model successfully identifies an image with a semantically relevant object and cultural descriptor(e.g., a Shinkansen). (ii) Object Relevant, Language-biased: The model successfully identifies an image with a semantically relevant object, but is distracted by the culture associated with the query language (e.g., a Thai train); (iii) Object Relevant: The model successfully identifies an image with a semantically relevant object, but the culture is irrelevant to the query language and the cultural descriptor (e.g., British Train). (iv) Cultural Descriptor Relevant: The model successfully identifies an image associated with the cultural descriptor, but misses the object’s semantics (e.g., Sushi). (v) Language-biased: The model fails to identify an image with a semantically relevant object but matches the query language’s culture (e.g., Thai dance). (vi) Totally Irrelevant: The model captures neither the relevant object nor the culture associated with the descriptor or language. (e.g., Eiffel Tower).
To quantify the "tugging effect" between implicit language bias and explicit instructions, we calculate the Similarity Drift ($\Delta\text{sim}$). Given a base concept query $q_{C}$ and a culturally-augmented query $q_{C+CD}$, the drift for an image $I$ is defined as:
$\Delta\text{sim}(I)=S_{q_{C+CD},I}-S_{q_{C},I}$ (5)
A higher $\Delta\text{sim}$ for the Correct category relative to other categories with mismatch culture indicates a model’s ability to prioritize explicit descriptors over inherent culture-language associations.